OpenStudy (compassionate):

I've look around OpenStudy and the net, but can't seem to find an understandable tutorial. "Find the diameter of a cone that has a volume of 83.74 cubic inches and a height of 5 inches. Use 3.14 for pi."

OpenStudy (anonymous):

Work backwards from the volume formula 1/3pir^2h

OpenStudy (anonymous):

\[\huge \frac{ 1 }{ 3 } (3.14)r^25 = 83.74\]

OpenStudy (compassionate):

V = 1/3 (p)(r^2)(h) > "Find the diameter of a cone that has a volume of 83.74 cubic inches and a height of 5 inches. Use 3.14 for pi." 83.74 = 1/3 (3.14)(r^2)(5)

OpenStudy (compassionate):

Would I have to isolate r?

OpenStudy (anonymous):

oh whoops

OpenStudy (anonymous):

multiply both sides by 3 and divide both sides by 3.14 and 5

OpenStudy (anonymous):

I got r^2 = 16.00127.. take the square root of both sides and you get: r = 4.0001

OpenStudy (anonymous):

basically 4 since we approximated pi

OpenStudy (compassionate):

83.74 = 1/3 (3.14)(r^2)(5) 251.22 = (3.14)(r^2)(5) 80.0063694268 = r^2 * 5 16.0012738854 = r^2 4.00015923251 = r

OpenStudy (anonymous):

correct. we can verify this by plugging the numbers back into the original equation

OpenStudy (compassionate):

However, 4.00015923251 is the radius @cleats , we're being asked to find the diameter. (radius is half the diameter) So we would say 2 x 4.00015923251 = 8.00031846502 Correct?

OpenStudy (anonymous):

oh, that part went over my head. Yes. Radius is half the diameter. Correct

OpenStudy (compassionate):

Thanks (:

OpenStudy (anonymous):

:)*

OpenStudy (compassionate):

You should just report the user.

OpenStudy (anonymous):

we did but the nazimods always seem to be around when I'm doing the naughty things

OpenStudy (anonymous):

anyhow, apologies for going off topic. Have fun bby

OpenStudy (compassionate):

>Nazimods Hehe. You're funny. Have a good day. (;